Calculate the pH at 25

°

C of a 0.13 M solution of a weak base with a Kb of 3.3

×

10−11.
Answer #1

Answer

pH = 8.32

Explanation

If the weak base sympolisd as B, the dissociation equation is

B(aq) + H2O(l)  <------> BH+(aq) + OH-(aq)

Kb = [BH+][OH-]/[B] = 3.3×10-11

Initial concentration

[B] = 0.13

[BH+] = 0

[OH-] = 0

Change in concentration

[B] = -x

[BH+] = +x

[OH-] = +x

Equillibrium concentration

[B] = 0.13 - x

[BH+] = x

[OH-] = x

x2/(0.13 -x) = 3.3×10-11

solving for x

x = 2.0712×10-6

[OH-] = 2.0712×10-6M

pOH = 5.68

pH = 14 - pOH

pH = 14 - 5.68 = 8.32